OpenStudy (anonymous):

Could someone help me simply this, trig functions are throwing me off. (written inside)

OpenStudy (anonymous):

\[\frac{ mv(2-\sin(\phi) }{ 2m*\cos(\arctan (\frac{ \cos (\phi) }{ 2 } ))}\]